# Vision Language Modeling of Content, Distortion and Appearance for Image Quality Assessment

### v2

@@ I Introduction
Although deep learning-based BIQA methods have significantly advanced state of the art, the key challenge of lacking large labeled training dataset still remains. Existing IQA databases are generally inadequate in size. A common strategy to overcoming this difficulty is to train a model using synthetic images with artificial distortion and then regress the model to a small-scale target BIQA database. However, image distortions in the real world are complex and it is very difficult to simulate all possible real world scenarios. Another approach is to use self-supervised learning (SSL) [4, 13, 50] to overcome the lack of sufficient training data, as SSL is able to leverage large amount of unlabeled data. A much more fundamental challenge in BIQA is that human perception of quality can vary with the content of the image [39], thus image quality is not only affected by distortion but is also closely tied to image content. Therefore content and distortion confound the image quality assessment problem. The fact that it is very difficult to distinguish genuine image features such as high frequency signals from distortion such as compression artifacts makes developing BIQA models even harder.
@@
Therefore, the challenge of developing BIQA models is how to encode the semantic content, distortion and appearance appropriately and relate them to the final perceived image quality. In this paper, we have developed a new BIQA model termed Self-supervision and Vision-Language supervision Image QUality Evaluator (SLIQUE) which models the semantic content, distortion and appearance together to assess an image’s visual quality. To enable the development of SLIQUE, we contribute a large Text Annotated Distortion, Appearance and Content (TADAC) image database. The TADAC database is the first of its kind specifically designed to exploit vision-language modeling for IQA research. The key contributions of this paper are as follows:
@@
  • We have developed a systematic method for constructing large text annotated image databases designed for exploiting vision-language modeling for image quality assessment and present the Text Annotated Distortion, Appearance and Content (TADAC) database containing over 1.6 million images annotated with texts about their semantic contents, distortion characteristics and appearance properties. We used existing labels or automatic image captioning to annotate the semantic content, designed a list of suitable textual phrases for describing the distortion characteristics, and developed automatic algorithms for computing the appearance properties and annotated these properties with suitable textual descriptions. The TADAC database is the first of its kind that is annotated with all three types of quality relevant texts to enable the learning of high level knowledge about all possible factors affecting image quality. TADAC has enabled the development of the first BIQA model (SLIQUE) that jointly models semantic content, distortion and appearance. We will make TADAC publicly available.
@@ II-A Blind Image Quality Assessment
Traditional BIQA methods mainly focus on designing hand-crafted features as the representation to predict the quality scores. Among them, models based on natural scene statistics (NSS) are widely studied,such as NIQE[25]. Another kind of techniques is based on visual codebooks.
Recently, the success of many deep learning-based computer vision tasks has inspired a number of BIQA models based on deep networks. However, lacking large labeled training dataset is one of the major challenges. One of the strategies is to use various approaches to augment the training datasets. PaQ-2-PiQ [49] introduces a large IQA dataset that includes images and patches to train IQA models. MUSIQ [17] uses multi-scale data for training. The model in [2] predicts the overall quality score of the image by combining scores of image patches, and the model in [62] adopts meta-learning to achieve the training. The DB-CNN model [52] uses synthetic and authentic distortion images to finetune a dual-stream pretrained network. The staircase structure model in [40] integrates feature from intermediate layers into final feature representation through mixed database training strategy. The model in [43] implements graph attention network to aggregate features decomposed by fully-connected layers. The model in [54] have developed a multi-task learning method where they apply language-supervised multi-task learning to assess image quality, recognize scenes, and identify distortions. By using textual descriptions of distorted images, they create image-text pairs to finetune the CLIP [32] model. GRepQ [38] learning low-level quality features agnostic to distortion types, Further, fine-tuning one such model to extract high-level image quality information through simple relevant text prompts.
@@ II-B Vision-Language Contrastive Learning
Vision-language contrastive learning is a mechanism that combines vision and language understanding. It learns to associate images and their textual descriptions in a shared embedding space. Through a contrastive loss function, the embedding textual vector and the visual vector of the same image are encouraged to be closer, while those of different images are pushed farther apart. Vision-language contrastive learning bridges the gap between vision and language, opening up new possibilities in image-text analysis. It is commonly used in multimodal tasks, including text-driven image manipulation [8, 47], image captioning [15], object detection [56, 37], and semantic segmentation [33, 58]. The majority of aforementioned works focus on understanding the high-level semantic of images through vision-language contrastive learning. The application of vision-language contrastive learning to learn image representations for BIQA has just started to attract the attention of researchers [44] where simple abstract textual descriptions like “good photo” and “bad photo” are used to fine-tune the pre-trained CLIP model [31]. However, as visual quality involves multiple influencing factors, it is insufficient to use simple abstract textual descriptions in providing enough textual information for image quality perception. Our method addresses the above limitations by designing and adopting text that is more diverse and related to image quality. We embeds text information into the model using a language supervision contrastive learning approach, enabling SLIQUE to more accurately extract image quality features.
@@ III SLIQUE
Our new BIQA model is termed Self-supervision and Vision-Language supervision Image QUality Evaluator (SLIQUE) and the training of its major part, the image encoder is shown in Fig. 2. The design of the SLIQUE architecture and training algorithm addresses the major challenges in BIQA as discussed previously. First, it employs self-supervised learning to exploit large amount of readily available unlabeled images. Second, it uses language-vision contrastive learning to embed high level quality relevant knowledge into the BIQA model. Third, it constructs a large Text Annotated Distortion, Appearance and Content (TADAC) image database for training the SLIQUE such that it can acquire quality relevant high level knowledge to accurately predict the visual quality of images.
@@ III-A Joint Contrastive Learning
$\mathcal{\textit{l}}=-\log\frac{\exp\left(q\cdot k_{+}/\tau\right)}{\sum_{i=0}^{N}\exp\left(q\cdot k_{i}/\tau\right)}$ (1)
@@
$\mathrm{\textit{L}}_{1}=\sum_{i}l\biggl{(}\ldots,F(x_{n}^{i}),G(t_{n}^{i}),\ldots\biggm{|}\left<x_{n}^{i},t_{n}^{i}\right>\in\mathcal{B}_{i},\\ 1\leq n\leq N\biggr{)}$ (2)
where $l(\cdot)$ is defined in Equation (1), $\mathcal{B}_{i}$ denotes the $i$-th training batch consisting of 1 positive sample pair and $N$-1 negative sample pairs, and $\langle x_{n}^{i},t_{n}^{i}\rangle$ is the $n$-th $(1\leq n\leq N)$ sample pair of image-language in $\mathcal{B}_{i}$. The pairing strategy can be found in Section III-D. Based on Equation (2), the image-language branch should be able to capture the intrinsic correlation between visual and textual representations.
@@
Another difference is that training samples prepared for this branch are image-image pairs rather than image-language ones. Thus, the constraint on the image-image branch can be expressed as Equation (3).
$\mathrm{\textit{L}}_{2}=\sum_{j}l\biggl{(}\ldots,F(y_{m}^{j}),F(z_{m}^{j}),\ldots\biggm{|}\left<y_{m}^{j},z_{m}^{j}\right>\in\mathcal{A}_{j},\\ 1\leq m\leq M\biggr{)}$ (3)
@@
$L_{joint}=({1-\alpha})L_{1}+\alpha L_{2}$ (4)
where $\alpha$ is the parameter to balance the image-image branch and the image-language branch. After minimizing Equation (4), the encoder parameters are frozen. The image encoder $F$ is used to extract image features which are then fed to a simple regression neural network to predict the image quality.
@@ III-B Regression to Quality Score
Similar to [1][22] , we use a simple ridge regression to map image features to a visual quality score. The regression is supervised by the Mean Opinion Score (MOS) from IQA datasets based on the $L_{1}$ loss. Note that given an image to be assessed, its textual description may be not comprehensive or even not available at all in real-world scenarios. Thus, the language encoder is excluded in the regression, and we do not require the textual description of an image to achieve the quality assessment in the testing.
@@ III-C Construction of database
In addition to synthetic distortion images, we collected 809,850 images with authentic (real) distortion from five datasets: 255,000 from AVA [28], 120,000 from COCO [21] training set, 33,000 from VOC [6], 400,000 from Places365 [57], and 1,850 from CERTH-Blur [23]. These datasets were originally built for various other purposes, e.g., the VOC database is introduced for the object recognition task. Thus, the labels for these images are also diverse. In brief, the images collected from CERTH-Blur are without any kinds of labels while the remaining images have scene category labels. The total number of category for these images is 26,332.
Content, Distortion and Appearance Annotation Texts: After collecting more than 1.6 million images, we need to prepare textual descriptions for them. As the joint contrastive learning is designed for the BIQA, the textual descriptions should be highly related to visual quality of images. It has been demonstrated that the visual quality of an image depends on its content, its distortion, and the interaction between them [60]. Thus, the annotation texts should describe the images content, distortion and their appearance characteristics.
For image content description text, we adopt the textual template format of “A photo of a(n) $\left\{s\right\}$”, where $\left\{s\right\}$ is given by the scene label or object label of the image. This template allows us to summarize the overall content information of the image based on $\left\{s\right\}$. Fortunately, most images collected above contain scene or object labels. For the images without semantic labels, we obtain their content description texts by employing the image captioning method in [26].
@@
For images containing authentic (real) distortions, they are much harder to describe textually because they may contain complex and multiple distortions. As the distortions are unknown, our approach is to annotate these images by their appearance characteristics based on 4 quality relevant aspects: brightness, contrast, sharpness, and colorfulness [42, 11]. For a given image, we first calculate its brightness (BR), contrast (CT), sharpness (SH), and colorfulness (CL) (detailed procedures for calculating these are described in Appendix B in the supplementary materials). We then classify these BR, CT, SH and CL values as “high”, “medium” and “low” categories based on their value distributions. For each value category within every aspect, we design 20 suitable descriptive phrases. In total we have $4\times 3\times 20=2400$ phrases for annotating images containing real distortions (see Appendix B in the supplementary material). For a given image, we label each aspect with a phrase randomly picked from the list corresponding to its value category. For example, for an image whose CT has a “medium” value, we randomly pick one of the 20 phrases corresponding to medium CT value. Note that in the current version of the TADAC datase, these appearance descriptions were assigned to the subset of images containing real distortions, they can be easily applied to the whole database.
@@ III-E Comparisons with Previous BIQA Works
As far we are aware, SLIQUE is the first BIQA model that explicitly models all three types visual quality relevant attributes, semantic contents, distortion characteristics and appearance properties. Our method is joint contrastive learning which combines vision-language supervised contrastive learning and visual self-supervised contrastive learning, and can take full advantage of these two contrastive learning methods. With the joint contrastive learning mechanism, our model is distinguished from the models that only use a vision-language supervised contrast learning mechanism (such as CLIP-IQA [44] and LIQE [54]), as well as the models with only visual self-supervised contrast learning (such as CONTRIQUE and Re-IQA). Removing either contrastive learning degrades the performance of the model as will be detailed in the ablation study of Table IV. Furthermore, we have constructed the first of a kind database specifically designed for learning all three categories of quality bearing image features. We have developed a systematic approaches to constructing quality relevant image textual labels for exploiting vision-language modeling technology to advance research in IQA.

### v3

@@ Abstract
© 2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
Although deep learning-based BIQA methods have significantly advanced state of the art, the key challenge of lacking large labeled training dataset still remains. Existing databases are generally inadequate in size. A common strategy to overcoming this difficulty is to train a model using synthetic images with artificial distortion and then regress the model to a small-scale target database. However, image distortions in the real world are complex and it is very difficult to simulate all possible real world scenarios. Another approach is to use self-supervised learning (SSL) [1, 2, 3] to overcome the lack of sufficient training data, as SSL is able to leverage large amount of unlabeled data. The proposed method [4] synthesizes perceptually aligned quality scores by fusing multiple Full-Reference IQA (FR-IQA) metrics, calibrating with human evaluations, and generating Synthetic Quality Benchmark (SQB) scores, enabling robust training of models for BIQA without extensive human labeling. A much more fundamental challenge in BIQA is that human perception of quality can vary with the content of the image [5, 6], thus image quality is not only affected by distortion but is also closely tied to image content. Therefore content and distortion confound the image quality assessment problem. The challenge in developing BIQA models lies in the difficulty of distinguishing genuine high-frequency image features from distortions caused by compression artifacts, as these can appear similar.
@@
Therefore, the challenge of developing BIQA models is how to encode the semantic content, distortion and appearance appropriately and relate them to the final perceived image quality. In this paper, we have developed a new model termed Self-supervision and Vision-Language supervision Image QUality Evaluator (SLIQUE) which models the semantic content, distortion and appearance together to assess an image’s visual quality. To enable the development of SLIQUE, we contribute a large Text Annotated Distortion, Appearance and Content (TADAC) image database. The database is the first of its kind specifically designed to exploit vision-language modeling for IQA research. The key contributions of this paper are as follows:
@@
  • We have developed a systematic method for constructing large text annotated image databases designed for exploiting vision-language modeling for image quality assessment and present the Text Annotated Distortion, Appearance and Content (TADAC) database containing over 1.6 million images annotated with texts about their semantic contents, distortion characteristics and appearance properties. We used existing labels or automatic image captioning to annotate the semantic content, designed a list of suitable textual phrases for describing the distortion characteristics, and developed automatic algorithms for computing the appearance properties and annotated these properties with suitable textual descriptions. The database is the first of its kind that is annotated with all three types of quality relevant texts to enable the learning of high level knowledge about all possible factors affecting image quality. TADAC has enabled the development of the first BIQA model (SLIQUE) that jointly models semantic content, distortion and appearance. We will make TADAC publicly available.
@@ II-A Blind Image Quality Assessment
Traditional methods mainly focus on designing hand-crafted features as the representation to predict the quality scores. Among them, models based on natural scene statistics (NSS) are widely studied,such as NIQE[7]. Another kind of techniques is based on visual codebooks.
Recently, the success of many deep learning-based computer vision tasks has inspired a number of BIQA models based on deep networks. However, lacking large labeled training dataset is one of the major challenges. One of the strategies is to use various approaches to augment the training datasets. PaQ-2-PiQ [8] introduces a large dataset that includes images and patches to train IQA models. MUSIQ [9] uses multi-scale data for training. The model in [10] predicts the overall quality score of the image by combining scores of image patches, and the model in [11] adopts meta-learning to achieve the training. The DB-CNN model [12] uses synthetic and authentic distortion images to finetune a dual-stream pretrained network. The staircase structure model in [13] integrates feature from intermediate layers into final feature representation through mixed database training strategy. The model in [14] implements graph attention network to aggregate features decomposed by fully-connected layers. The model in [15] have developed a multi-task learning method where they apply language-supervised multi-task learning to assess image quality, recognize scenes, and identify distortions. By using textual descriptions of distorted images, they create image-text pairs to finetune the CLIP [16] model. GRepQ [17] learning low-level quality features agnostic to distortion types, Further, fine-tuning one such model to extract high-level image quality information through simple relevant text prompts.
@@ II-B Vision-Language Contrastive Learning
Vision-language contrastive learning is a mechanism that combines vision and language understanding. It learns to associate images and their textual descriptions in a shared embedding space. Through a contrastive loss function, the embedding textual vector and the visual vector of the same image are encouraged to be closer, while those of different images are pushed farther apart. Vision-language contrastive learning bridges the gap between vision and language, opening up new possibilities in image-text analysis. It is commonly used in multimodal tasks, including text-driven image manipulation [18, 19], image captioning [20], object detection [21, 22], and semantic segmentation [23, 24]. The methods in[25, 26] introduce a triple contrastive learning framework for vision-language pre-training, leveraging both image-text and image-image to enhance cross-model representation learning. The majority of aforementioned works focus on understanding the high-level semantic of images through vision-language contrastive learning. The application of vision-language contrastive learning to learn image representations for BIQA has just started to attract the attention of researchers [27] where simple abstract textual descriptions like “good photo” and “bad photo” are used to fine-tune the pre-trained CLIP model [28]. However, as visual quality involves multiple influencing factors, it is insufficient to use simple abstract textual descriptions in providing enough textual information for image quality perception. Our method addresses the above limitations by designing and adopting text that is more diverse and related to image quality. We embeds text information into the model using a language supervision contrastive learning approach, enabling SLIQUE to more accurately extract image quality features.
@@ III SLIQUE
Our new model is termed Self-supervision and Vision-Language supervision Image QUality Evaluator (SLIQUE) and the training of its major part, the image encoder is shown in Fig. 2. The design of the SLIQUE architecture and training algorithm addresses the major challenges in BIQA as discussed previously. First, it employs self-supervised learning to exploit large amount of readily available unlabeled images. Second, it uses language-vision contrastive learning to embed high level quality relevant knowledge into the model. Third, it constructs a large Text Annotated Distortion, Appearance and Content (TADAC) image database for training the SLIQUE such that it can acquire quality relevant high level knowledge to accurately predict the visual quality of images.
@@ III-A Joint Contrastive Learning
$\mathcal{\textit{l}}=-\log\frac{\exp\left(q\cdot k_{+}/\tau\right)}{\sum_{i=0}^{N}\exp\left(q\cdot k_{i}/\tau\right)},$ (1)
@@
$\mathrm{\textit{L}}_{1}=\sum_{i}l\biggl{(}\ldots,F(x_{n}^{i}),G(t_{n}^{i}),\ldots\biggm{|}\left<x_{n}^{i},t_{n}^{i}\right>\in\mathcal{B}_{i},\\ 1\leq n\leq N\biggr{)},$ (2)
where $l(\cdot)$ is defined in Equation (1), $\mathcal{B}_{i}$ denotes the $i$-th training batch consisting of 1 positive sample pair and $N$-1 negative sample pairs, and $\langle x_{n}^{i},t_{n}^{i}\rangle$ is the $n$-th $(1\leq n\leq N)$ sample pair of image-language in $\mathcal{B}_{i}$.
The pairing strategy can be found in Section III-D. Based on Equation (2), the image-language branch should be able to capture the intrinsic correlation between visual and textual representations.
@@
Another difference is that training samples prepared for this branch are image-image pairs rather than image-language ones. The training objective of image-image branch is to enhances the model’s ability to capture subtle visual details that textual descriptions may miss. The constraint on the image-image branch can be expressed as Equation (3).
$\mathrm{\textit{L}}_{2}=\sum_{j}l\biggl{(}\ldots,F(y_{m}^{j}),F(z_{m}^{j}),\ldots\biggm{|}\left<y_{m}^{j},z_{m}^{j}\right>\in\mathcal{A}_{j},\\ 1\leq m\leq M\biggr{)},$ (3)
@@
$L_{joint}=({1-\alpha})L_{1}+\alpha L_{2},$ (4)
where $\alpha$ is the parameter to balance the image-image branch and the image-language branch. Notably, $L_{1}$ and $L_{2}$ have the same value range, as both of them are based on the InfoNCE loss. Thus, the role of the parameter $\alpha$ is to weight two terms rather than scaling either of them. After minimizing Equation (4), the encoder parameters are frozen. The image encoder $F$ is used to extract image features which are then fed to a simple regression neural network to predict the image quality.
@@ III-B Regression to Quality Score
Following the suggestion in [31], we employ the ridge regression, a regularized linear regressor, to map the encoded image feature h to a visual quality score $y$:
${y=\mathbf{w}\cdot\mathbf{h},}$ (5)
where $w$ is the regression vector estimated by
${\mathbf{w}=\underset{\mathbf{w}}{\text{argmin}}\sum_{k}(GT_{k}-y_{k})^{2}+\lambda\\ |\mathbf{w}\|^{2},}$ (6)
where $GT_{k}$ and $y_{k}$ are the MOS and the predicted score of the $k$-th image in the training set respectiveily, and $\lambda$ is a regularization constant ($\lambda$ is empirically set to 0.2 in this work). Note that given an image to be assessed, its textual description may be not comprehensive or even not available at all in real-world scenarios. Thus, the language encoder is excluded in the regression, and we do not require the textual description of an image to achieve the quality assessment in the testing.
@@ III-C Construction of database
In addition to synthetic distortion images, we collected 809,850 images with authentic (real) distortion from five datasets: 255,000 from AVA [40], 120,000 from COCO [41] training set, 33,000 from VOC [42], 400,000 from Places365 [43], and 1,850 from CERTH-Blur [44]. These datasets were originally built for various other purposes, e.g., the AVA dataset is for aesthetic visual analysis, VOC and COCO datasets are introduced for object recognition task. Thus, the labels for these images are also diverse. In brief, the images collected from CERTH-Blur are without any kinds of labels while the remaining images have scene category labels. The total number of category for these images is 26,332.
The dataset compilation process results in a well-balanced distribution of synthetic and authentic distortion images. Specifically, we acquire 840,000 synthetically distorted images and 809,850 authentically distorted images, yielding an approximate 1:1 ratio between the two categories. This balanced composition enhances the dataset’s utility for training and evaluating machine learning models in image distortion analysis.
Content, Distortion and Appearance Annotation Texts: After collecting more than 1.6 million images, we need to prepare textual descriptions for them. As the joint contrastive learning is designed for the BIQA, the textual descriptions should be highly related to visual quality of images. It has been demonstrated that the visual quality of an image depends on its content, its distortion, and the interaction between them [30]. The semantic content of an image pertains to its high-level information, encompassing elements such as objects, scenes, and contextual relationships. In contrast, both distortions and appearance are concerned with low-level image quality attributes. Specifically, distortion represents explicit aspects of image quality, whereas appearance embodies implicit characteristics that influence perceptual assessment. Thus, the annotation texts should describe the images content, distortion and their appearance characteristics.
For image content description text, we adopt the textual template format of “A photo of a(n) $\left\{s\right\}$”, where $\left\{s\right\}$ is given by the scene label or object label of the image. This template allows us to summarize the overall content information of the image based on $\left\{s\right\}$. Fortunately, most images collected above contain scene or object labels. For the images without semantic labels, we obtain their content description texts by employing the image captioning method in [45]. To ensure the quality and consistency of the caption, we apply length-based filtering to eliminate excessively short or long captions. Subsequently, we conduct a manual review to verify the accuracy and relevance of the captions.
@@
(a) Brightness
(b) Contrast
(c) Sharpness
(d) Colorfulness
Figure 3: The distribution of four quality relevant aspects of authetic image in TADAC (a)Brightness, (b)Contrast, (c)Sharpness and (d)Colorfulness
For images containing authentic (real) distortions, they are much harder to describe textually because they may contain complex and multiple distortions. As the distortions are unknown, our approach is to annotate these images by their appearance characteristics based on 4 quality relevant aspects: brightness, contrast, sharpness, and colorfulness [46, 47]. For a given image, we first calculate its brightness (BR), contrast (CT), sharpness (SH), and colorfulness (CL) (detailed procedures for calculating these are described in Appendix B in the supplementary materials). The distribution of the four appearance attribute values are shown in Fig. 3(d). Analysis revealed a distribution skewed towards median values, with fewer extremes. To balance the dataset, we manually established thresholds for BR, CT, SH, and CL values, categorizing them as “high,” “medium,” or “low.” This manual classification scheme is calibrated to ensure an approximately equal number of images across the three categories. Recent studies have demonstrated the efficacy of single-modal GPT [48] to be a reliable evaluation instrument for purely linguistic tasks [49].
For each level within each visual aspect, we generate a set of descriptive phrases, such as “$\left\{l\right\}$ contrast,” “The image is $\left\{l\right\}$ contrast” or “The contrast in image is $\left\{l\right\}$”, where $l$ is the level of the contrast. While these phrases may vary lexically, they maintain semantic consistency. In total we have $4\times 3\times 20=240$ phrases for annotating images containing real distortions (see Appendix B in the supplementary material). For a given image, we label each aspect with a phrase randomly picked from the list corresponding to its value category. For example, for an image whose CT has a “medium” value, we randomly pick one of the 20 phrases corresponding to medium CT value. Note that in the current version of the TADAC dataset. These appearance descriptions were assigned to the subset of images containing real distortions, they can be easily applied to the whole database.
@@ III-E Comparisons with Previous BIQA Works
As far we are aware, SLIQUE is the first BIQA model that explicitly models all three types visual quality relevant attributes, semantic contents, distortion characteristics and appearance properties. Our method is joint contrastive learning which combines vision-language supervised contrastive learning and visual self-supervised contrastive learning, and can take full advantage of these two contrastive learning methods. With the joint contrastive learning mechanism, our model is distinguished from the models that only use a vision-language supervised contrast learning mechanism (such as CLIP-IQA [27] and LIQE [15]), as well as the models with only visual self-supervised contrast learning (such as CONTRIQUE and Re-IQA). Removing either contrastive learning degrades the performance of the model as will be detailed in the ablation study of Table VI. Furthermore, we have constructed the first of a kind database specifically designed for learning all three categories of quality bearing image features. We have developed a systematic approach to constructing quality relevant image textual labels for exploiting vision-language modeling technology to advance research.
@@ IV-A Implementation Details
TABLE III: Performance comparisons of SLIQUE against various BIQA models on underwater IQA databases.
Method UIF UID2021 Weighted Average
[…]
HyperIQA 0.290 0.311 0.476 0.450 0.383 0.381
MUSIQ 0.309 0.285 0.403 0.383 0.281 0.265
CONTRIQUE 0.704 0.731 0.833 0.830 0.770 0.781
CLIP-IQA+ 0.425 0.429 0.298 0.328 0.251 0.267
SLIQUE 0.766 0.792 0.860 0.853 0.814 0.823
TABLE IV: Performance comparisons of SLIQUE against various BIQA models on enhanced IQA databases.
Method QADS GFIQA Average
[…]
HyperIQA 0.747 0.710 0.880 0.892 0.814 0.801
MUSIQ 0.772 0.768 0.910 0.901 0.841 0.835
CONTRIQUE 0.935 0.938 0.911 0.914 0.923 0.926
CLIP-IQA+ 0.675 0.680 0.870 0.877 0.773 0.779
SLIQUE 0.959 0.953 0.932 0.930 0.946 0.942